### v1

@@ 1 Introduction
Increasing polarization in U.S. politics has sparked interest in the diverging economic implications of each of the major political parties’ policy platforms. For instance, recent Democratic presidential candidates have pushed for investment in the clean energy sector and a reduction in CO2 emissions, exemplified by policies like the Inflation Reduction Act of 2022, while Republican candidates have supported the increased extraction of fossil fuels through expanded drilling and deregulation. Moreover, a large body of empirical literature demonstrates that news about changes in future fundamentals affects the economy today by altering expectations (for a review, see Beaudry and Portier, 2014). However, little work has been done to study the potential economic effects of news about future policy that is transmitted through election events.
@@
As an alternative to the baseline and narrative election shocks, I also asses the response of stock prices to a measure that ignores election probabilities and only considers election outcomes. The response of clean energy prices remains qualitatively similar to that of the baseline measure, while the responses of energy and defense stocks display no persistent effects. These differences highlight the limitations of the alternative measure, which fails to capture anticipatory effects before election day and the degree of uncertainty resolved by each election outcome.
@@
Related Literature. The findings of this paper are closely linked with the literature that takes the timing of fiscal policy shocks seriously. Notably, Ramey (2011) shows that much of the response to defense spending increases happen in anticipation of policy implementation. Romer and Romer (2010) argue that expectations of tax changes play little role in the response of the macroeconomy relative to their actual implementation. Leeper et al. (2012, 2013) map news of government spending and tax changes into the DSGE framework and argue that accounting for time-variation in the news process is vital. Ricco et al. (2016) argues that clear policy communications sharpen the response to fiscal spending through the consolidation of market expectations. Most recently, and perhaps most related to this paper, Hazell and Hobler (2024) exploit high-frequency variation in asset prices immediately after the 2020 Georgia Senate runoff to estimate the effects of a budget deficit shock. Relative to these papers I provide two key insights. First, I show that not just election outcomes, but also changes in election probabilities can affect stock markets through updates to investors expectations about future government policy. Second, I shift the focus from aggregate outcomes to sectoral reallocation by showing that elections are an important source of fluctuations in sectors that are differentially affected by the industrial policy platforms of each political party.
@@
My work also relates to the literature studying the financial and economic effects of political uncertainty. A key insight in this literature is that higher political uncertainty leads to higher stock market volatility (Pástor and Veronesi, 2013; Kelly et al., 2016; Goodell et al., 2020; Mnrasi and Essaddam, 2021). Relative to this work, I shift the focus from stock market volatility to stock market returns and directly link the policy platforms of the political parties to sector-specific performance. The labor market application in this paper also closely relates to Hassan et al. (2019) who find that firms react to increased political risk by reducing hiring and investment. In contrast with my findings, their political risk measure identifies little variation at the sectoral level. In other words, at least in the energy, clean energy, and defense sectors, the response of stock prices and employment to election shocks suggests that firm-level political risk is correlated.
@@
Finally, from a methodological standpoint, this paper broadly contributes to the high-frequency time series identification literature. The pioneering work in this area has largely focused on identifying monetary policy shocks (Gürkaynak et al., 2005; Gertler and Karadi, 2015; Nakamura and Steinsson, 2018; Miranda-Agrippino and Ricco, 2021; Bauer and Swanson, 2023a, 2023b; Bianchi et al., 2023). Recent work has extended these methods to study oil supply shocks (Kanzïg, 2021; Degasperi, 2024; Polat et al., 2025), uncertainty shocks (Ferrara and Guérin, 2018), and Treasury supply shocks (Phillot, 2025). Closer in spirit to this paper is Dreschel (2024) who studies the effects of “political pressure” shocks to the Federal Reserve on macroeconomic conditions. To my knowledge, no existing work has utilized election betting data to identify election shocks.
@@
The remainder of the paper proceeds as follows. Section 2 discusses “politicized sectors,” i.e., sectors that are particularly affected by the different policy platforms of the two major U.S. political parties and provides some preliminary evidence that the stock markets of these sectors react to election news. Section 3 delineates the construction of the election shock series and evaluates it by studying its movements near narrative events. Section 4 presents the effects of the election shocks on financial and labor markets. Section 5 concludes.
@@ 2.1 Industrial Policy and Sectoral Partisanship
Although the particular policy platforms of the Republican and Democratic parties have evolved over the past several decades, and vary substantially at the candidate-level, the industrial policy stances of each party relative to the other have remained consistent in several aspects. I opt to focus on three sectors: energy, clean energy, [^1] Throughout, energy refers to the “traditional” energy sector comprised mostly of oil and coal production, while clean energy refers to the production of energy by alternative sources that do no produce emissions, i.e., nuclear, hydroelectric, solar, wind, and geothermal energy. and defense, which throughout my sample (2000-2024), have been differentially affected by the two parties. I claim that, relative to the Democratic party, the Republican party’s proposed and implemented policies propagate the expansion of the energy and defense sectors and contraction of the clean energy sector. In this section I provide narrative and empirical evidence in support of this claim.
@@
Starting with the energy and clean energy sectors, first note that in many cases policies that encourage production of the former disfavor the latter and vice-versa. To illustrate, consider the implementation of carbon taxes, which lowers the demand for coal and oil while increasing the demand for clean energy sources (Aldy, 2013; Hafstead and Picciano, 2017; Macalauso et al., 2018). This directly follows from the fact that carbon taxes increase the cost of consuming coal and oil, and clean energy sources are substitutes for these traditional energy sources. Similarly, Acemoglu et al. (2012) argue that interventions in “dirty” sectors leads to more technological innovation in “clean” sectors. However, recent impactful bills may affect the production and technological innovation of both sources of energy in the same direction, e.g., the Inflation Reduction Act of 2022 (IRA) created tax credits clean energy generation while extending and increasing credits for industrial carbon capture. [^2] Carbon capture are technologies used to capture C02 emissions from industrial power plants. Many large oil companies like Exxon Mobil and Shell have recently invested in these technologies. Despite these caveats, for the majority of legislation discussed in this section there is a clear winner between the two sectors.
@@
Table A1 provides a comprehensive list of significant legislation passed by recent Republican and Democratic administrations that directly affect the energy or clean energy sector. In each case a brief description of the relevant policies are included. Some notable Republican policies are the Bush administration’s tax credits for refined coal production under the American Jobs Creation Act, the Trump administration’s reversal of permit pipeline denials and vehicle C02 emissions standards, and withdrawal from the Paris Climate Agreement (PCA) in 2017 and 2025. Note that, while the PCA does not contain any formal constraints, it legally binds participating countries to share progress on reducing emissions, and features a commitment from developed countries to provide financial and technological assistance to developing countries in reaching its goals. Thus, a country entering or withdrawing from the PCA largely serves as a signal about its subsequent energy approach.
@@
Significant Democratic legislation has conversely leaned towards favoring the clean energy sector and disfavoring the energy sector, at least relative to Republican administrations. Several bills have inhibited production of coal and oil, such as the Obama administration’s banning of new coal mining leases, and the Biden administration’s raising of royalties to drill on federal lands. Each administration also passed notable legislation intended to directly increase clean energy production. For instance, Obama’s American Recovery and Reinvestment Act and Biden’s IRA have contributed an estimated \$90 billion (\$265 billion) in clean energy investment and tax incentives. Finally, the Obama and Biden administrations approved the construction of fewer new gas pipelines than the Bush and Trump administrations. In particular, from 2001-2024 Republican administrations approved an average of 180% more miles of pipe for major pipeline projects per year than Democratic administrations. Details are plotted in Figure A1.
In some instances Republican (Democratic) administrations have passed pro-clean energy (energy) policies. Notably, Bush’s 2005 tax credits for solar energy investment and nuclear power generation; and Biden’s 2022 extension and increase of tax credits for carbon capture. However, even in these cases the policies are often only an auxiliary provision of an otherwise large bill with more significant investments in the opposing sector. For example, Biden’s carbon capture tax credits were a provision in the IRA, which brought much larger investments to clean energy as discussed above. Furthermore, while we cannot predict what an administration of the opposing party would have passed under the same circumstances, in some instances there is evidence that a similar bill would have had more aggressive policies benefiting the party’s favored sector. For example, after taking office in 2008 Obama removed a \$2,000 cap on the solar energy tax credits signed by Bush, i.e., an Energy Policy Act of 2005 signed by Obama may have had more pro-clean energy provisions than the one in our history books. [^3] Interestingly, Obama voted in favor of the bill while his eventual 2008 primary and general election opponents Hillary Clinton (Democrat) and John McCain (Republican) voted against it. With these caveats in mind, it is important to clarify that for the later analysis in Section 4 I simply require that firms in each sector seek to gain more from one of the two parties.
@@
Figure 1 plots U.S. defense spending as a percentage of GDP from 1993 to 2024 fiscal years (FY). [^4] Fiscal years in the federal government start in October 1st of the preceding year and end September 30th, e.g., FY2022 starts October 1, 2021 and ends September 30, 2022. Vertical lines indicate the first fiscal year that the defense budget was set by a particular administration. Notably, from start to end of term, defense spending increased under each Republican administration and decreased under each Democratic administration. More specifically, the budget increased substantially from 3.9% to 5.4% from FY2001 to FY2009 under the Bush administration, by FY2017 had decreased to 3.8% under the Obama administration, increased modestly to 4% by FY2021 under the Trump administration, and finally decreased modestly again under the Biden administration to 3.6% by FY2024. The large increase in defense spending during the Bush administration was at least partially driven by the wars in Afghanistan and Iraq, while declines during the Obama administration are linked with the significant withdrawal of U.S. troops from these countries. It is difficult to disentangle the exogeneity of these events, e.g., would the Iraq War have taken place under alternative leadership? However, there is significant evidence provided by a collection of speeches from this period as well as the political science literature that indicates under an administration led by Al Gore, Bush’s 2000 general election opponent, the war would not have taken place (see for example Gore, 2002; Harvey, 2012). [^5] Similarly, during the 2008 presidential elections Obama’s general election opponent, John McCain criticized Obama’s proposed 16 month deadline for withdrawal of troops in Iraq and offered a longer timeline (for a brief overview see Cooper et al., 2008).
@@
Aside from studying passed legislation and budgets, to study which sectors of the economy seek to gain more from one of the two political parties one can observe their lobbying activities. To the extent that firms’ donations are driven by the proposed policies of candidates as opposed to other factors (e.g., social values), these donations reveal their monetary preferences. Figure 2 plots the contributions from the energy, alternative energy, and defense sectors to each political party during the 2000-2024 presidential elections. On aggregate, energy firms consistently donate the most between the three sectors, though there is significant heterogeneity between election years. There is a clear transition between 2008 and 2012 where Republican contributions nearly doubled from \$60 million to \$105 million. This discontinuity is likely driven by Citizens United v. Federal Election Commission (2010), which unrestricted independent expenditures of corporations for political campaigns. Nonetheless, firms in the energy sector donated more than twice as much to Republican candidates compared with Democratic candidates in each election, indicating a clear preference. [^6] The energy sector also includes clean energy firms, though they make up a small enough proportion contributors that it does not significantly bias contributions to Democratic candidates upwards. The alternative energy sector, which includes clean energy and bioenergy firms, displays bias in the opposite direction, albeit with substantially lower aggregate contributions. Aside from 2012, firms in this category donated more than twice as much to Democratic candidates as Republican candidates. Note that the contributions of bioenergy companies lean Republican [^7] For instance, of the top 10 alternative energy contributors in the 2024 election cycle, 2 were bioenergy companies and each donated significantly more to Republican candidates than Democratic candidates. and thus the bias in donations towards the Democratic party in the alternative energy serves as a lower bound for that of the clean energy sector. Finally, firms in the defense sector favored Republican candidates in all but the 2008 election, with approximately 34% more donations to the Republican party on average. The milder differences in contributions in this sector may be partially driven by the fact that defense companies heavily rely on contracts with the federal government, i.e., it is advantageous for these companies to lobby with both parties. Indeed, in the 2024 election the top 5 defense contributors each donated at least \$500,000 to both political party.
@@ 2.2 Election Events and Stock Markets
Having established the relative political preferences of these sectors, a question of interest is whether sectoral stock markets react to election events. Events that provide news about the probability that one of the two political parties will gain political power also provide news about future industrial policy. To illustrate, consider the release of negative information about “Candidate A,” which reduces her probability of winning the election and increases the probability of “Candidate B.” Suppose that Candidate B has proposed increasing taxes on the production of oil, while Candidate A denounces this proposal. Then the increase to Candidate B’s probability can be viewed as a negative shock to the expectation of future oil production, the profits of oil producing firms, and even the supply of other goods produced by firms that use oil as an input in their production process. All else equal, this innovation also serves as a positive shock to the expectation of future demand for oil substitutes, e.g., nuclear energy. There is substantial evidence that stock markets react to signals about future monetary policy (Swanson, 2021; Del Negro et al., 2023), fiscal policy (Baker et al., 2019), and firm-level political risk (Hassan et al., 2019). This suggests that, to the extent that investors are able to observe and interpret election news as containing signals about future industrial policy, the asset prices of affected sectors will respond.
@@
Figure 3 provides preliminary narrative evidence of this mechanism. Each bar plots the percentage change in the aggregate asset price of a given sector between the day before and the day after a significant election event. Prices are given by the relevant S&P subindices. [^8] The Bloomberg tickers for energy, clean energy, and defense are SPN, SPGTCED, and SPSIAD, respectively. I consider three election events from the 2024 election: the June 27th debate between Joe Biden and Donald Trump, the September 10th debate between Kamala Harris and Donald Trump, and Donald Trump’s presumptive election victory on November 5th. The first of these events provided a negative signal about Biden’s chances at reelection, and was concurrent with a moderate decrease of approximately 2.4% in clean energy prices. The Harris debate, for which Harris was the consensus winner, improved the Democrats’ chances at the White House and energy and clean energy prices changed by -2.9% and 2.9%, respectively. Finally, Trump’s victory on election day is associated with substantial contemporaneous increases to energy and defense prices of 4% and 5%, and a decrease in clean energy prices of 6.5%.
@@
These price movements are suggestive, but a more formal analysis is required to understand the extent to which stock markets respond contemporaneously to innovations in election probabilities. It is possible that these movements are particular to the 2024 election, perhaps due to the intricacies of Biden, Harris, or Trump’s policy platforms. Furthermore, it is important to consider whether any stock market effects are persistent. Kwon and Tang (2023) argue that investor beliefs overreact (underreact) to extreme (subtle) types of corporate news leading to reversal (drift) in asset prices. Thus the short-term reactions suggested by Figure 3 may only capture the overexcitement of investors. Finally, does election news have any real effects? If firms update their expectations the same way that investors do, they may alter their hiring or production decisions before any bills are passed, or even before the inauguration of the next president. The stock market performance of a firm may also indirectly affect its output, e.g., after an increase to asset prices, financially constrained firms are more likely to raise capital through the equity-financing channel (Baker et al., 2003; Campello and Graham, 2013). In the next section I construct an election shock series to study these possibilities.

### v2

@@ 1 Introduction
Increasing polarization in U.S. politics has sparked interest in the diverging economic implications of each of the major political parties’ policy platforms. For instance, recent Democratic presidential candidates have pushed for investment in the clean energy sector and a reduction in CO2 emissions, exemplified by policies like the Inflation Reduction Act (IRA) of 2022, while Republican candidates have supported the increased extraction of fossil fuels through expanded drilling and deregulation. Moreover, a large body of empirical literature demonstrates that news about changes in future fundamentals affects the economy today by altering expectations (for a review, see Beaudry and Portier, 2014). However, little work has been done to study the potential economic effects of news about future policy that is transmitted through election events.
@@
As an alternative to the baseline and narrative election shocks, I also asses the response of stock prices to a measure that ignores election probabilities and only considers election outcomes. The response of clean energy prices remains qualitatively similar to that of the baseline measure, while the responses of energy and defense stocks are not persistent. These differences highlight the limitations of the alternative measure, which fails to capture anticipatory effects before election day and the degree of uncertainty resolved by each election outcome.
@@
Related Literature. The findings of this paper are closely linked with the literature that takes the timing of fiscal policy shocks seriously. Notably, Ramey (2011) shows that much of the response to defense spending increases happen in anticipation of policy implementation. Romer and Romer (2010) argue that expectations of tax changes play little role in the response of the macroeconomy relative to their actual implementation. Leeper, Richter and Walker (2012); Leeper, Walker and Yang (2013) map news of government spending and tax changes into the DSGE framework and argue that accounting for time-variation in the news process is vital. Ricco, Callegari and Cimadomo (2016) argues that clear policy communications sharpen the response to fiscal spending through the consolidation of market expectations. Most recently, and perhaps most related to this paper, Hazell and Hobler (2024) exploit high-frequency variation in asset prices immediately after the 2020 Georgia Senate runoff to estimate the effects of a budget deficit shock. Relative to these papers I provide two key insights. First, I show that not just election outcomes, but also changes in election probabilities can affect stock markets through updates to investors expectations about future government policy. Second, I shift the focus from aggregate outcomes to sectoral reallocation by showing that elections are an important source of fluctuations in sectors that are differentially affected by the industrial policy platforms of each political party.
@@
My work also relates to the literature studying the financial and economic effects of political uncertainty. A key insight in this literature is that higher political uncertainty leads to higher stock market volatility (Pástor and Veronesi, 2013; Kelly, Pástor and Veronesi, 2016; Goodell, McGee and McGroarty, 2020). Relative to this work, I shift the focus from stock market volatility to stock market returns and directly link the policy platforms of the political parties to sector-specific performance. The labor market application in this paper also closely relates to Hassan et al. (2019) who find that firms react to increased political risk by reducing hiring and investment. In contrast with my findings, their political risk measure identifies little variation at the sectoral level.
@@
Finally, from a methodological standpoint, this paper broadly contributes to the high-frequency time series identification literature. The pioneering work in this area has largely focused on identifying monetary policy shocks (Gürkaynak, Sack and Swanson, 2005; Gertler and Karadi, 2015; Nakamura and Steinsson, 2018; Miranda-Agrippino and Ricco, 2021; Bauer and Swanson, 2023a, b; Bianchi et al., 2023). Recent work has extended these methods to study oil supply shocks (Känzig, 2021; Degasperi, 2023; Polat et al., 2025), uncertainty shocks (Ferrara and Guérin, 2018), and Treasury supply shocks (Phillot, 2025). Closer in spirit to this paper is Drechsel (2024) who studies the effects of “political pressure” shocks to the Federal Reserve on macroeconomic conditions. To my knowledge, no existing work has utilized election betting data to identify election shocks.
@@
The remainder of the paper proceeds as follows. Section 2 discusses “politicized sectors,” i.e., sectors that are particularly affected by the different policy platforms of the two major U.S. political parties, and provides some preliminary evidence that the stock markets of these sectors react to election news. Section 3 delineates the construction of the election shock series and evaluates it by studying its movements near narrative events. Section 4 presents the effects of the election shocks on financial and labor markets. Section 5 concludes.
@@ 2.1 Industrial Policy and Sectoral Partisanship
Although the particular policy platforms of the Republican and Democratic parties have evolved over the past several decades, and vary substantially at the candidate-level, the industrial policy stances of each party relative to the other have remained consistent in several aspects. I opt to focus on three sectors — energy, [^1] Throughout, energy refers to the “traditional” energy sector comprised mostly of oil and coal production, while clean energy refers to the production of energy by alternative sources that do no produce emissions, i.e., nuclear, hydroelectric, solar, wind, and geothermal energy. clean energy, and defense — which, throughout my sample period (2000–2024), have been differentially affected by the two parties. I claim that, relative to the Democratic party, the Republican party’s proposed and implemented policies to propagate the expansion of the energy and defense sectors and contraction of the clean energy sector. In this section I provide narrative and empirical evidence in support of this claim.
@@
Starting with the energy and clean energy sectors, first note that in many cases policies that encourage production of the former disfavor the latter and vice-versa. To illustrate, consider the implementation of carbon taxes, which lowers the demand for coal and oil while increasing the demand for clean energy sources (Aldy, 2013; Hafstead and Picciano, 2017; Macaluso et al., 2018). This directly follows from the fact that carbon taxes increase the cost of consuming coal and oil, and clean energy sources are substitutes for these traditional energy sources. Similarly, Acemoglu et al. (2012) argue that intervention in “dirty” sectors leads to more technological innovation in “clean” sectors. However, recent impactful bills may affect the production and technological innovation of both sources of energy in the same direction, e.g., the IRA created tax credits clean energy generation while extending and increasing credits for industrial carbon capture. [^2] Carbon capture are technologies used to capture C02 emissions from industrial power plants. Many large oil companies like Exxon Mobil and Shell have recently invested in these technologies. Despite these caveats, for the majority of legislation discussed in this section one of the two sectors is a clear relative beneficiary.
@@
Table LABEL:taba1 provides a comprehensive list of significant legislation passed by recent Republican and Democratic administrations that directly affected the energy or clean energy sector. In each case a brief description of the relevant policies are included. Some notable Republican policies are the Bush administration’s tax credits for refined coal production under the American Jobs Creation Act, the Trump administration’s reversal of permit pipeline denials and vehicle C02 emissions standards, and Trump’s withdrawal from the Paris Climate Agreement (PCA) in 2017 and 2025. Note that, while the PCA does not contain any formal constraints, it legally binds participating countries to share progress on reducing emissions, and features a commitment from developed countries to provide financial and technological assistance to developing countries. Thus, a country entering or withdrawing from the PCA serves as a signal about its subsequent energy approach.
@@
Significant Democratic legislation has conversely leaned towards favoring the clean energy sector and disfavoring the energy sector, at least relative to Republican administrations. Several bills have inhibited production of coal and oil, such as the Obama administration’s banning of new coal mining leases, and the Biden administration’s raising of royalties to drill on federal lands. Each administration also passed notable legislation intended to directly increase clean energy production. For instance, Obama’s American Recovery and Reinvestment Act and Biden’s IRA have contributed an estimated \$90 billion (\$265 billion) in clean energy investment and tax incentives. Finally, the Obama and Biden administrations approved the construction of fewer new gas pipelines than the Bush and Trump administrations. In particular, from 2001-2024 Republican administrations approved an average of 180 percent more miles of pipe for major pipeline projects per year than Democratic administrations. Details are plotted in Figure A1.
Republican (Democratic) administrations have at times enacted policies favoring clean energy (energy), such as Bush’s 2005 tax credits for solar and nuclear investment and Biden’s 2022 expansion of carbon capture incentives. However, these measures often appear as secondary provisions within broader bills that heavily invest in the opposite sector—for example, Biden’s carbon capture credits were part of the IRA, which primarily supported clean energy. While we cannot know how the opposing party would have legislated under similar conditions, some evidence suggests they may have pursued stronger policies in their preferred sector. For instance, upon taking office, Obama lifted a \$2,000 cap on Bush’s solar tax credit—suggesting an Energy Policy Act of 2005 signed by Obama might have leaned more pro-clean energy. With these caveats in mind, the analysis in Section 4 assumes only that firms in each sector expect to benefit more under one party than the other.
@@
Figure 1 plots U.S. defense spending as a percentage of GDP from 1993 to 2024 fiscal years (FY). [^3] Fiscal years in the federal government start in October 1st of the preceding year and end September 30th, e.g., FY2022 starts October 1, 2021 and ends September 30, 2022. Vertical lines indicate the first fiscal year that the defense budget was set by a particular administration. Notably, from start to end of term, defense spending increased under each Republican administration and decreased under each Democratic administration. More specifically, the budget increased substantially from 3.9 to 5.4 percent from FY2001 to FY2009 under the Bush administration, by FY2017 had decreased to 3.8 percent under the Obama administration, increased modestly to 4 percent by FY2021 under the Trump administration, and finally decreased modestly again under the Biden administration to 3.6 percent by FY2024. The large increase in defense spending during the Bush administration was at least partially driven by the wars in Afghanistan and Iraq, while declines during the Obama administration are linked with the significant withdrawal of U.S. troops from these countries. It is difficult to disentangle the exogeneity of these events, e.g., would the Iraq War have taken place under alternative leadership? However, there is significant evidence provided by a collection of speeches from this period as well as the political science literature that indicates under an administration led by Al Gore, Bush’s 2000 general election opponent, the war would not have taken place (see for example Gore, 2002; Harvey, 2012). [^4] Similarly, during the 2008 presidential elections Obama’s general election opponent, John McCain criticized Obama’s proposed 16 month deadline for withdrawal of troops in Iraq and offered a longer timeline (for a brief overview see Cooper et al., 2008).
@@
Aside from studying passed legislation and budgets, to study which sectors of the economy seek to gain more from one of the two political parties one can observe their lobbying activities. To the extent that firms’ donations are driven by the proposed policies of candidates as opposed to other factors (e.g., social values), these donations reveal their monetary preferences. Figure 2 plots the contributions from the energy, alternative energy, and defense sectors to each political party during the 2000-2024 presidential elections. On aggregate, energy firms consistently donate the most between the three sectors, though there is significant heterogeneity between election years. There is a clear transition between 2008 and 2012 where Republican contributions nearly doubled from \$60 million to \$105 million. This discontinuity is likely driven by Citizens United v. Federal Election Commission (2010), which unrestricted independent expenditures of corporations for political campaigns. Nonetheless, firms in the energy sector donated more than twice as much to Republican candidates compared with Democratic candidates in each election, indicating a clear preference. [^5] The energy sector also includes clean energy firms, though they make up a small enough proportion contributors that it does not significantly bias contributions to Democratic candidates upwards. The alternative energy sector, which includes clean energy and bioenergy firms, displays bias in the opposite direction, albeit with substantially lower aggregate contributions. Aside from 2012, firms in this category donated more than twice as much to Democratic candidates as Republican candidates. Note that the contributions of bioenergy companies lean Republican [^6] For instance, of the top 10 alternative energy contributors in the 2024 election cycle, 2 were bioenergy companies and each donated significantly more to Republican candidates than Democratic candidates. and thus the bias in donations towards the Democratic party in the alternative energy serves as a lower bound for that of the clean energy sector. Finally, firms in the defense sector favored Republican candidates in all but the 2008 election, with approximately 34 percent more donations to the Republican party on average. The milder differences in contributions in this sector may be partially driven by the fact that defense companies heavily rely on contracts with the federal government, i.e., it is advantageous for these companies to lobby with both parties. Indeed, in the 2024 election the top 5 defense contributors each donated at least \$500,000 to both political party.
@@ 2.2 Election Events and Stock Markets
Having established the relative political preferences of these sectors, a question of interest is whether sectoral stock markets react to election events. Events that provide news about the probability that one of the two political parties will gain political power also provide news about future industrial policy. To illustrate, consider the release of negative information about “Candidate A,” which reduces her probability of winning the election and increases the probability of “Candidate B.” Suppose that Candidate B has proposed increasing taxes on the production of oil, while Candidate A denounces this proposal. Then the increase to Candidate B’s probability can be viewed as a negative shock to the expectation of future oil production, the profits of oil producing firms, and even the supply of other goods produced by firms that use oil as an input in their production process. All else equal, this innovation also serves as a positive shock to the expectation of future demand for oil substitutes, e.g., nuclear energy. There is substantial evidence that stock markets react to signals about future monetary policy (Swanson, 2021; Del Negro, Giannoni and Patterson, 2023), fiscal policy (Baker et al., 2019), and firm-level political risk (Hassan et al., 2019). This suggests that, to the extent that investors observe and interpret election news as containing signals about future industrial policy, the asset prices of affected sectors will respond.
@@
Figure 3 provides preliminary narrative evidence of this mechanism. Each bar plots the percentage change in the aggregate asset price of a given sector between the day before and the day after a significant election event. Prices are given by the relevant S&P subindices. [^7] The Bloomberg tickers for energy, clean energy, and defense are SPN, SPGTCED, and SPSIAD, respectively. I consider three election events from the 2024 election: the June 27th debate between Joe Biden and Donald Trump, the September 10th debate between Kamala Harris and Donald Trump, and Donald Trump’s presumptive election victory on November 5th. The first of these events provided a negative signal about Biden’s chances at reelection, and was concurrent with a moderate decrease of approximately 2.4 percent in clean energy prices. The Harris debate, for which Harris was the consensus winner, improved the Democrats’ chances at the White House and energy and clean energy prices changed by -2.9 and 2.9 percent, respectively. Finally, Trump’s victory on election day is associated with substantial contemporaneous increases to energy and defense prices of 4 and 5 percent, and a decrease in clean energy prices of 6.5 percent.
@@
These price movements are suggestive, but a more formal analysis is required to understand the extent to which stock markets respond contemporaneously to innovations in election probabilities. It is possible that these movements are particular to the 2024 election, perhaps due to the intricacies of Biden, Harris, or Trump’s policy platforms. Furthermore, it is important to consider whether any stock market effects are persistent. Kwon and Tang (2025) argue that investor beliefs overreact (underreact) to extreme (subtle) types of corporate news leading to reversal (drift) in asset prices. Thus the short-term reactions suggested by Figure 3 may only capture the overexcitement of investors. Finally, does election news have any real effects? If firms update their expectations the same way that investors do, they may alter their hiring or production decisions before any bills are passed, or even before the inauguration of the next president. The stock market performance of a firm may also indirectly affect its output, e.g., after an increase to asset prices, financially constrained firms are more likely to raise capital through the equity-financing channel (Baker, Stein and Wurgler, 2003; Campello and Graham, 2013). In the next section I construct an election shock series to study these possibilities.